1. Find the general solution of the differential equation.
y′′′−6y′′−36y′+216y=0. Use C1, C2, C3, ... for the constants of integration.

y(t) = _______

2. Determine a suitable form for Y(t) if the method of undetermined coefficients is to be used.
Do not evaluate the constants.

y′′′−4y′ = te^−2t + 3cos(2t)

Use J,K,L,M as coefficients. Enclose arguments of functions in parentheses. For example, sin(2x). Do not simplify trigonometric functions of nt, where n is a positive integer.
1
Expert's answer
2020-07-07T19:20:25-0400

Solution:

1.y′′′−6y′′−36y′+216y=0


k3-6k2-36k+216=0

k2(k-6)-36(k-6)=0

(k-6)(k2-36)=0

(k-6)2(k+6)=0

k1=k2=6,k3=-6

y(t)=C1e6t+C2te6t+C3e-6t


2.y′′′−4y′ = te-2t + 3cos(2t)


k3-4k=0

k(k2-4)=0

k(k-2)(k+2)=0

k1=-2, k2=0, k3=2

y0(t)=C1e-2t+C2+C3e2t

yp1(t)=(Jt2+Kt)e-2t

yp2(t)=Lcos(2t)+Msin(2t)

y(t)=C1e-2t+C2+C3e2t+(Jt2+Kt)e-2t+Lcos(2t)+Msin(2t)

Answer:

1.y(t)=C1e6t+C2te6t+C3e-6t

2.y(t)=C1e-2t+C2+C3e2t+(Jt2+Kt)e-2t+Lcos(2t)+Msin(2t)